# High Frequency Quoting Under Liquidity Constraints

### v1

@@ 10 Reference Leg Selection via Order Book Liquidity Profiles
The two approaches discussed in this paper reflect different assumptions about the information content of market data. The Hawkes-based rule treats reference-impacting events as informative flow variables, modeling their interactions over time. In contrast, the CLF-based rule assumes that execution-relevant information is embedded in the instantaneous structure of the limit order book. This distinction is not meant to imply any claim about informational efficiency; rather, it frames a practical and philosophical contrast between flow-based and state-based decision systems for short-term execution modeling.
11 Experiment
We now evaluate the reference leg selection rules introduced earlier using high-frequency limit order book data from February 7, 2022, for NIFTY futures contracts expiring in February ($F_{c}$) and March ($F_{n}$). Each method—Hawkes-based forecasting, CLF-based slope analysis, and the market-optimal benchmark—is assessed within the unified decision framework described in Problem 1.2.
@@ 11 Experiment
Hawkes-Based Forecasting. The decision rule $\chi^{\mathcal{H}}(w)$, defined in Section 9, uses simulated forward intensities derived from a fitted multivariate Hawkes process. For each rolling window of 1000 ms, we estimate the Hawkes kernels over reference-impacting events—trades, cancellations, and downward price modifications—using maximum likelihood. We then simulate $K$ future realizations over a 10 ms forecast horizon. These simulations yield reference-side event counts, from which arrival ratios $\rho^{X,Y}_{(w)}$ are computed and used to derive the binary decision $\chi^{\mathcal{H}}(w)\in\{0,1\}$.
CLF-Based Rule. In parallel, we implement the order-book-state-based decision rule $\chi^{\mathcal{B}}(w)$, described in Section 10. At each tick, we compute CLF scores over multiple depths (denoted $CLF_{1}$ through $CLF_{4}$) for both $F_{c}$ and $F_{n}$. A lower CLF value indicates a more favorable liquidity slope for execution. The contract with the lower CLF is selected at each tick, and a majority vote over the previous 1000 ms determines the decision $\chi^{\mathcal{B}}(w)$ used for the next interval. Ties are resolved in favor of the next-month contract $F_{n}$.
Modeling Remark Although directional liquidity values used in the CLF-based rule are computed at tick resolution, the formal decision framework in Problem 1.2 requires that all quoting rules be evaluated over fixed intervals $w\in\mathcal{W}$. To reconcile this, the book-based filter $\mathcal{G}^{\mathcal{B}}(w)\in\mathbb{R}^{2}$ is computed by applying an exponential moving average (EMA) to tick-level CLF values observed over the historical window $(w_{1}-h,w_{1}]$. The resulting smoothed features are assumed to remain valid throughout the forecast interval $w=[w_{1},w_{2})$, and are used to produce a deterministic reference leg decision $\chi^{\mathcal{B}}(w)$.
@@
Market-Optimal Benchmark. Prior to computing the Hawkes based rule and the CLF based rule, we use the method presented in Section 7 to compute a benchmark for evaluation of both rules. We construct a reference oracle $\chi^{m}(w)\in\{0,1\}$, which selects the contract that would have yielded the lower realized spread in hindsight. For each window $w\in\mathcal{W}$, we compute the actual execution spreads for both reference leg choices and select the contract that minimizes slippage.
@@
Evaluation Metric. Each method’s prediction accuracy is measured using the agreement score $\mathcal{A}(\chi,\chi^{m})$, as defined in Problem 1.2:
@@
This metric reflects how often each rule’s decision matches the market-optimal choice based on realized execution outcomes.
Decision Rule Summary. To contextualize the empirical comparison, we summarize the three decision rules evaluated in this experiment.
@@
12 Results
We divide the results into two parts. First, we evaluate the relative predictive performance of different Hawkes kernel specifications using the test for superior predictive ability (SPA). Second, we compare the predictive accuracy of the Hawkes-based reference leg selection rule $\chi^{\mathcal{H}}(w)$ and the CLF-based rule $\chi^{\mathcal{B}}(w)$, both benchmarked against the market-optimal decision $\chi^{m}(w)$ introduced in Problem 1.2.
@@ 12.1 Hawkes Kernel Selection
Table 8 reports SPA test values for four kernel specifications: two parametric (HawkesExp and HawkesSumExp) and two non-parametric (HawkesEM and HawkesCondLaw). The parametric methods yield comparable test values of 0.017 and 0.015, respectively, indicating similar performance. The non-parametric methods exhibit greater variability: the EM-based kernel records a substantially higher test value of 0.530, while the conditional law kernel yields 0.0. These results suggest that the expectation-maximization approach offers superior predictive ability under this test.
@@
Since the SPA test does not reject the null hypothesis that the EM-based kernel is the benchmark, we adopt the HawkesEM specification for all subsequent analyses.
@@ 12.2 Comparison of Hawkes Arrival Ratio and CLF
We now compare the Hawkes-based rule $\chi^{\mathcal{H}}(w)$ and the CLF-based rule $\chi^{\mathcal{B}}(w)$ against the market-optimal benchmark $\chi^{m}(w)$, using the agreement score $\mathcal{A}(\chi,\chi^{m})$ defined in Problem 1.2. The evaluation uses high-frequency tick data from February 7, 2022, for NIFTY futures contracts $F_{c}$ and $F_{n}$.
For the Hawkes rule, we fit the expectation-maximization kernel over a rolling 1000 ms window, advanced in 10 ms steps, using only events that impact top-of-book price (trades, cancellations, downward price moves). For each forecast window of 10 ms, we simulate arrival counts and compute arrival ratios $\rho^{X,Y}_{(w)}$, selecting the reference leg via the decision rule defined in Section 9.
The CLF-based rule is evaluated using four variants $CLF_{1}$ through $CLF_{4}$, corresponding to increasing quote depth. At each tick, the contract with the lower CLF value is selected. These raw tick-level decisions are smoothed via an exponential moving average over the interval $(\tau-h,\tau]$, as described in the Experiment section, and the resulting signal is used to produce a binary decision $\chi^{\mathcal{B}}(w)$ for each forecast window.
The market-optimal benchmark $\chi^{m}(w)$ is computed using realized spread over each minute-long interval. The agreement score $\mathcal{A}(\chi,\chi^{m})$ is then used to measure the fraction of windows for which each method’s predicted reference leg matches the benchmark.
Table 9: Agreement score $\mathcal{A}(\chi,\chi^{m})$: Proportion of windows where predicted reference leg matches market-optimal decision (Feb 7, 2022)
@@
Hawkes Arrival Ratio 0.7335
The results in Table 9 show that the Hawkes arrival ratio significantly outperforms all CLF variants, correctly identifying the optimal reference leg in 73.35% of windows. Among the CLF variants, performance improves with quote depth, with $CLF_{4}$ achieving 63.58%, followed by $CLF_{3}$, $CLF_{2}$, and $CLF_{1}$.
These findings highlight the predictive advantage of the Hawkes framework, which models temporal cross-excitation in event arrivals. By capturing dependencies between price-impacting events across contracts and book sides, the Hawkes method provides a dynamic signal for price stability, whereas CLF scores rely on static order book snapshots. This result supports the inclusion of event dynamics in real-time reference leg selection for spread execution strategies.

### v2

@@ 3.6 Reference Leg Selection via LOB Liquidity Profiles
The two approaches discussed in this paper reflect different assumptions about the information content of market data. The Hawkes-based rule treats reference-impacting events as informative flow variables, modeling their interactions over time. In contrast, the CLF-based rule assumes that execution-relevant information is embedded in the instantaneous structure of the limit order book.
4 Data
This section describes the tick-driven market data and the microstructural features used by our filters. We use tick-by-tick (TBT) data from the National Stock Exchange (NSE) of India; the experiment analyzes NIFTY futures on February 7, 2022.333All exchange specifications and market microstructure rules referenced in this study reflect those in effect as of March 2022, corresponding to the period from which the data were collected. The dataset includes two instruments: the current-month contract expiring on February 24, 2022 ($F_{c}$), and the next-month contract expiring on March 31, 2022 ($F_{n}$). As shown in Table 6, both contracts display substantial tick activity on the sample date, providing granularity for execution modeling.
Each tick is one of four event types: NEW (order addition), MODIFY (order revision), CANCEL (order withdrawal), and TRADE (execution). These types are used to extract features relevant to quoting behavior and limit order book dynamics. The minimum tick size for NIFTY futures is —0.05, as mandated by NSE.
In our methodology, this data is segmented into a collection of evaluation windows $\mathcal{W}=\{w=[w_{1},w_{2})\}$, where each window spans a short time interval (e.g., 100 milliseconds) aligned with the quoting decision frequency. For each $w\in\mathcal{W}$, we apply two distinct classes of filters:
  1. The tick-based filter $\mathcal{G}^{\mathcal{T}}(w)$ operates on the order flow observed strictly before the quoting time $w_{1}$.
  2. The LOB-based filter $\mathcal{G}^{\mathcal{B}}(w)$ extracts features from the limit order book snapshot observed exactly at time $w_{1}$.
Tick-derived features.
For each evaluation window $w\in\mathcal{W}$, the tick-based filter $\mathcal{G}^{\mathcal{T}}(w)$ encodes summary statistics of the recent order flow observed strictly before the quoting time $w_{1}$. The features chosen as filtration criteria are those event types that worsen the price stability of the selected reference contract. These features include:
  1. Count of executed trades (TRADE events).
  2. Count of order cancellations (CANCEL events).
  3. Count of price-worsening quote updates (MODIFY events where the best bid decreases or the best ask increases).
  4. Total number of tick events, inclusive of all above types.
LOB-derived features.
Analogous to tick-derived features, we define features derived from the current limit order book (LOB) to evaluate reference-leg selection. In each interval $w\in\mathcal{W}$, the LOB-based filter $\mathcal{G}^{\mathcal{B}}(w)$ extracts directional liquidity indicators from the limit order book snapshot observed at time $w_{1}$. These features are computed separately for each contract leg and for each relevant side of the LOB (bid or ask), based on whether the leg is expected to be lifted or hit under the quoting convention.
Specifically, we extract:
  1. Price at each level $p^{b,i},p^{a,i}$, for $i=1,\dots,\nu$
  2. Quantity available at each level $q^{b,i},q^{a,i}$, for $i=1,\dots,\nu$
These level-wise features are later used to construct directional liquidity measures such as the Composite Liquidity Factor (CLF), introduced in a subsequent section, which serve as inputs to the LOB-based quoting rule $\mathcal{Q}^{\mathcal{B}}$.
5 Experiment and Results
From a market microstructure perspective, the two decision rules are designed to respond to different aspects of market information. The CLF is computed directly from limit order book primitives: instantaneous depth, queue imbalance, and local LOB shape, which are widely used as proxies for short-horizon liquidity and price impact in limit-order markets [2, 17]. This construction makes the CLF-based rule sensitive to rapid changes in the relative liquidity of the two legs over short horizons. However, during episodes of intense trading activity, high-frequency quotes and trades are heavily affected by microstructure noise and fleeting orders. Hasbrouck and Saar [27] document that low-latency trading can amplify such transitory LOB dynamics. In such environments, an exclusively cross-sectional, LOB-based statistic can generate spurious or very short-lived signals.
The Hawkes arrival-ratio rule, in contrast, is a low-dimensional summary of the event-time dynamics of price-impacting messages. It aggregates information about how often price-impacting events arrive on each leg, with an explicit memory structure. Consistent with empirical evidence that Hawkes-based event-time models capture clustering and persistent order-flow dynamics [4, 26, 15], we expect this construction to be relatively insensitive to purely transitory fluctuations in the order-flow and better suited to detecting and exploiting persistent regimes in the benchmark.
We deliberately differentiate the information sets of the two rules. For the Hawkes arrival-ratio we use only reference-impacting events at the top of the book (arrivals, cancellations, and downward price modifications that change the best quotes). For the CLF-based rule, we construct LOB-state measures from the first five depth levels on each side, thereby capturing the local cross-sectional shape of the LOB. In this sense, the comparison pits a history of reference-impacting ticks against a smoothed history of LOB snapshots. Our empirical analysis focuses on February 7, 2022 and on NIFTY futures expiring in February ($F_{c}$) and March ($F_{n}$). Throughout, the Hawkes arrival-ratio rule and the CLF-based LOB-state rule are evaluated within the decision-function framework of Section 2, using the oracle $\chi^{m}(w)$ as a benchmark. Table 6 reports tick-event counts for each contract on February 7, 2022. Counts include all four TBT event types (NEW, MODIFY, CANCEL, TRADE). The substantial volume supports evaluation of quoting policies at high temporal resolution.
Contract Date Tick events (count)
NIFTY February Futures 2022-02-07 49,828,303
NIFTY March Futures 2022-02-07 16,349,197
Table 6: Tick-event counts by contract on February 7, 2022. Counts include NEW, MODIFY, CANCEL, and TRADE events. A sample from the processed tick data is provided in the appendix, in Table A2 for illustration.
The evaluation period contains both short-lived bursts in which the oracle reference leg flips frequently and extended segments in which the oracle remains on the same leg for many consecutive windows. The presence of both low- and high-persistence benchmark regimes within a single day is useful for our purposes: it allows us to examine how order-flow based and LOB-based rules behave under different benchmark regimes while keeping the market context fixed. Among the runs, Figure 4 shows that benchmark regimes are distributed almost symmetrically across the two legs: across the four bins $Q_{1}$$Q_{4}$ the oracle generates 124 runs on $F_{c}$ and 123 runs on $F_{n}$. This symmetry emphasises that benchmark regimes are not dominated by a single contract in terms of the number of episodes. In the subsequent analysis we investigate this structure by conditioning agreement scores on the run-length and oracle’s benchmark contract, so that the relative performance of the Hawkes arrival-ratio and CLF can be interpreted in terms of both the duration and the contract choice of benchmark regimes.
[…]
Figure 4: Number of oracle runs in each benchmark-persistence bin on February 7, 2022, split by the benchmark’s selected reference leg.
[…]
Figure 5: Share of evaluation windows in each benchmark-persistence bin on February 7, 2022.
To study how performance varies with the stability of the oracle, we construct what we call benchmark persistence. Starting from the sequence of oracle decisions $\{\chi^{m}(w)\}_{w}$, we segment the evaluation period into maximal contiguous runs during which the oracle selects the same reference leg. The length of a run, measured in evaluation windows, is our measure of persistence: very short runs correspond to rapid benchmark switches between contracts, whereas long runs correspond to extended periods in which the oracle consistently prefers the same leg.
We then examine the empirical distribution of run lengths on February 7, 2022 and partition runs into bins of benchmark persistence. The first bin contains runs of length 1–3 windows, the second bin contains runs of length 4–10 windows, the third bin collects runs of length 11–58 windows, and the fourth bin consists of runs of length at least 59 windows. Figures 4 and 5 summarise this structure. Figure 4 reports the number of oracle runs in each persistence bin, while Figure 5 shows the share of evaluation windows that fall into each bin. The first two bins account for many distinct runs (72 and 52, respectively) but together cover only about $1.3\%$ of evaluation windows. The third bin contains 62 medium-length runs and accounts for roughly $4.6\%$ of windows. The fourth bin also contains 62 runs, but these long, high-persistence regimes account for about $94\%$ of the evaluation period.
Splitting runs by the oracle’s preferred contract shows that benchmark persistence is driven more by regime duration than by regime count. Across all four bins $Q_{1}$$Q_{4}$, the numbers of runs on the two contracts are similar. Over the full evaluation period, however, the oracle selects the near leg in about $72\%$ of windows. Within the highest-persistence bin $Q_{4}$, runs on the near leg last on average roughly three times as long as runs on the other leg and together account for close to $69\%$ of all windows. By contrast, the low- and medium-persistence bins $Q_{1}$$Q_{3}$ cover only about $6\%$ of the day and are more evenly split across the two contracts. Benchmark regimes are therefore concentrated in long intervals during which one leg remains market-optimal, punctuated by shorter excursions in which the oracle switches to the other leg.
5.1 Hawkes-Based Forecasting.
@@ 5 Experiment and Results
Within each decision interval $w$, the rule $\chi^{\mathcal{H}}(w)$ in Section 3.4 uses simulated forward intensities from a multivariate Hawkes model. Using only data available at the start of $w$, we estimate the Hawkes kernels over the past 1000 ms for the reference-impacting events by maximum likelihood, then simulate $K$ forward paths over the next 10 ms. From the simulated reference-side event counts, we compute arrival ratios $\rho^{X,Y}_{(w)}$ and apply the prescribed threshold to issue the binary reference-leg decision $\chi^{\mathcal{H}}(w)\in\{0,1\}$.
5.2 CLF-Based Rule.
At each tick, we compute CLF at depths 1–4 for both $F_{c}$ and $F_{n}$ $(CLF_{1},\dots,CLF_{4})$. CLF contrasts the cross-level price gap with cumulative depth, so that a lower value indicates more favourable immediate liquidity on the consumed side. The contract with the lower per-tick CLF score is selected at that tick as the reference contract. We partition the evaluation period into decision windows $w$ of length $10\,\text{ms}$, shifted forward in $10\,\text{ms}$ steps. For each window $w$, we take a majority vote over the tick-level reference choices observed in the preceding $1{,}000\,\text{ms}$. This yields the reference-leg decision $\chi^{\mathcal{B}}(w)$ applied to the next $10\,\text{ms}$ interval, with ties broken in favour of the next-month contract $F_{n}$.
Modeling Remark Tick-level CLF values must be mapped to the interval-based decision framework of Section 2. We construct the LOB-based filter $\mathcal{G}^{\mathcal{B}}$ by applying an exponential moving average (EMA) to per-tick CLF observations over the historical window $(w-\Delta,\,w)$, using only information available before the start of the interval. Let $c_{t}$ denote the CLF at tick time $t<w$. In our implementation we set $\Delta=1{,}000\,\text{ms}$, so that the EMA smooths CLF over the same 1,000 ms history used in the majority-vote construction and in the Hawkes-based rule’s estimation window.
5.3 Market-Optimal Benchmark.
@@
Before computing the Hawkes-based rule and the CLF-based rule, we use the method presented in Section 3.2 for evaluation. We construct a reference oracle $\chi^{m}(w)\in\{0,1\}$, which selects the contract that would have yielded the lower realized spread in hindsight. For each window $w\in\mathcal{W}$, we compute the actual execution spreads for both reference leg choices and select the contract that minimizes slippage.
5.4 Evaluation Metric.
@@
Each method’s prediction accuracy is measured using the agreement score $\mathcal{A}(\chi,\chi^{m})$, as defined in Section 2:
@@
This metric reflects how often each rule’s decision matches the market-optimal choice based on realized execution outcomes. To contextualize the empirical comparison, we summarize the three decision rules evaluated in this experiment.
@@
We divide the results into two parts. First, we compare alternative Hawkes kernel specifications using the test for superior predictive ability [25]. Second, we evaluate the reference-leg selection rules, $\chi^{\mathcal{H}}(w)$ and $\chi^{\mathcal{B}}(w)$, using deviation from the target spread, with both benchmarked against the oracle $\chi^{m}(w)$ introduced in Section 2.
@@ 5.5 Hawkes Kernel Selection
Table 7 reports SPA [25] test values for four kernel specifications: two parametric (HawkesExp and HawkesSumExp) and two non-parametric (HawkesEM and HawkesCondLaw). The parametric methods yield comparable test values of 0.017 and 0.015, respectively, indicating similar performance. The non-parametric methods exhibit greater variability: the EM-based kernel records a substantially higher test value of 0.530, while the conditional law kernel yields 0.0. The results suggest that the expectation-maximization kernel [33] is not outperformed by any of the other candidate models when used as a benchmark.
@@
SPA setup and interpretation.
For each candidate specification $k$, we treat $k$ as the benchmark and test whether any competing model achieves a lower expected loss than $k$ under the chosen loss function. We report the SPA p-value $p_{k}$ for this benchmarked comparison. Small values of $p_{k}$ lead to rejection of the null (there is evidence that at least one competing model outperforms the benchmark $k$); large values of $p_{k}$ imply we fail to reject the null (no evidence, in this sample, that any competitor is superior to benchmark $k$). Accordingly, higher p-values do not establish optimality, but they indicate that the chosen benchmark withstands the SPA comparison on this sample [25].
Since the SPA test does not reject the null hypothesis that the EM-based kernel is the benchmark, we adopt the HawkesEM specification for all subsequent analyses. Consistent with evidence that high-frequency order-flow intensities exhibit time-of-day patterns [39], we report intraday kernel variation, as shown in Figures 8.
@@ 5.6 Comparison of Hawkes Arrival Ratio and CLF
Using the forecasting rules defined in the preceding subsections and the agreement metric $\mathcal{A}(\chi,\chi^{m})$ introduced in Section 2, we compare the Hawkes-based rule $\chi^{\mathcal{H}}(w)$ and the CLF-based rule $\chi^{\mathcal{B}}(w)$ against the market-optimal oracle $\chi^{m}(w)$. The evaluation uses high-frequency tick data from February 7, 2022 for the NIFTY futures contracts $F_{c}$ and $F_{n}$. All CLF parameters and Hawkes kernel coefficients are estimated on a training sample; the evaluation period is used only for generating forecasts and for computing agreement scores. The oracle benchmark $\chi^{m}(w)$ is the only object that uses realised post-decision information when defining the market-optimal reference leg.
Table 8: Agreement score $\mathcal{A}(\chi,\chi^{m})$: proportion of evaluation windows in which each method’s predicted reference leg matches the market-optimal oracle on February 7, 2022. The joint-agreement row reports the fraction of windows in which both the Hawkes arrival ratio and CLF4 simultaneously match the oracle.
@@
Hawkes arrival ratio 0.6856
Hawkes–CLF4 joint agreement 0.4149
Table 8 shows that, on this evaluation day, the Hawkes arrival-ratio rule attains an agreement score of $0.6856$ with the market-optimal oracle. Among the CLF variants, CLF4 achieves the highest agreement score at $0.6358$, followed by CLF3, CLF2, and CLF1. On February 7, 2022, therefore, the low-dimensional Hawkes arrival ratio extracted from the limit-order dynamics attains a modestly higher agreement with the oracle than the best-performing CLF specification.
To quantify how often the two rules arrive at the same decision, Figure 6 and the joint-agreement row of Table 8 report the fraction of evaluation windows in which each combination of outcomes occurs. In $41.49\%$ of windows, both the Hawkes arrival ratio and CLF4 match the oracle. The Hawkes arrival ratio is uniquely in agreement with the oracle in $27.08\%$ of windows, while CLF4 is uniquely in agreement in $22.10\%$. Only $9.34\%$ of windows are misclassified by both rules simultaneously.
Conditional on the Hawkes arrival ratio agreeing with the oracle, CLF4 is also in agreement in about $60.5\%$ of windows; conditional on CLF4 agreeing with the oracle, the Hawkes arrival ratio is also in agreement in about $65.2\%$. These conditional shares indicate substantial overlap between the two rules, together with a non-trivial set of windows in which only one of them matches the oracle. A hypothetical aggregator that could choose, ex ante, the better of the two signals in each window would miss the benchmark leg in at most the $9.34\%$ of windows where both rules disagree with the oracle.
Panels (b) and (c) of Figure 6 separate agreement by the oracle’s selected contract for each rule in isolation. In these panels, the rows distinguish evaluation windows in which the oracle selects the current-month contract $F_{c}$ from those in which it selects the next-month contract $F_{n}$, while the columns indicate whether the rule matches or deviates from the oracle. Panel (b) shows that the Hawkes arrival-ratio rule agrees with the oracle in approximately $88\%$ of windows when the oracle selects $F_{c}$, but only in about $18\%$ of windows when the oracle selects $F_{n}$. Panel (c) shows that CLF4 attains agreement of about $61\%$ when the oracle selects $F_{c}$ and about $69\%$ when it selects $F_{n}$. Taken together, these contract-wise panels indicate that the Hawkes arrival ratio is most informative in regimes where the current-month contract is the market-optimal reference leg, whereas CLF4 contributes relatively more information when the oracle favours the next-month contract.
[…]
Figure 6: Joint agreement of the Hawkes arrival-ratio rule and CLF4 relative to the oracle on February 7, 2022. Panel (a) reports unconditional shares of evaluation windows in each joint outcome. Panel (b) reports contract-wise agreement and disagreement of the Hawkes arrival-ratio rule with the oracle. Panel (c) reports contract-wise agreement and disagreement of CLF4 with the oracle. Each cell reports the percentage and the number of evaluation windows corresponding to that outcome, using a common colour palette across panels.
Given the partition in 4, each evaluation window inherits the benchmark-persistence bin of the run to which it belongs. Figure 7 reports agreement scores for the Hawkes arrival-ratio and CLF4 within each persistence bin. In the lowest-persistence bin (runs of length 1–3 windows), CLF4 attains an agreement score of approximately $0.459$, compared with $0.406$ for the Hawkes arrival ratio. In the second bin (runs of length 4–10 windows), CLF4 reaches an agreement score of $0.524$ and the Hawkes arrival ratio reaches $0.489$. In the third bin (runs of length 11–58 windows), CLF4 attains an agreement score of $0.576$, while the Hawkes arrival ratio attains $0.413$. In the highest-persistence bin (runs of at least 59 windows), the ordering reverses: the Hawkes arrival-ratio rule reaches $0.702$ agreement with the oracle, while CLF4 attains $0.641$.
[…]
(a) Agreement by benchmark-persistence bin
[…]
(b) Difference in agreement (Hawkes - CLF4) by bin
Figure 7: Agreement scores of the Hawkes arrival-ratio and CLF4 across benchmark-persistence bins. Panel (a) shows agreement relative to the oracle in each bin of benchmark persistence. Panel (b) displays the corresponding difference in agreement (Hawkes minus CLF4); CLF4 attains higher agreement in the three lower-persistence bins (negative bars), while the Hawkes arrival ratio attains higher agreement in the highest-persistence bin (positive bar).
The aggregate agreement scores, the benchmark-persistence and contract-wise decompositions describe a consistent pattern. The Hawkes arrival-ratio rule and the LOB-based CLF4 rule both achieve non-trivial agreement with the market-optimal oracle, with a large intersection of windows in which they select the same reference leg, yet they extract information from different parts of the joint distribution of order-flow and LOB state. The persistence analysis shows that most evaluation windows fall into long benchmark regimes, within which the arrival-ratio rule is comparatively more accurate, while the CLF4 rule performs relatively better in short and medium-lived regimes in which the oracle switches reference legs more frequently. The contract-wise panels further indicate that the arrival ratio is most informative when the benchmark favours the current-month contract $F_{c}$, whereas CLF4 contributes more when the benchmark favours the next-month contract $F_{n}$. These findings suggest that event-history and LOB-state information are complementary rather than interchangeable inputs for reference-leg selection. Figures 4 and 6 summarise this structure, including the contract-wise split of runs across the two contracts. They motivate the state-dependent interpretation of the two complementary signals developed in the discussion that follows.